Please do not omit any steps.
Use the method of Undetermined Coefficients to solve y – 2y = 8 sin’a; y(0) = 0,7(0) = 0.
Answer #1

1 - 4 cosau. m - 2m = 0 y - 2y! = 8 sch²x. y - 2y = 84 ½ - 1 cos2x) 9-2y = 4 Auxillary equation of mim- 2) = 0 complementa& y(x) = Ye(x) typ (x) 22 4x) - cit cze 2a + schaat cola 2 y (0%=0 citoz ta [C, + c ₂ = - 1 / 1 y(x) = gege 2x 2 + cos2x - s